# Determinant and Pfaffian formulas for particle annihilation

### v1

@@ 5.3. The coalescence-to-annihilation map
The key observation is that pairwise coalescence can be reinterpreted as complete annihilation via a parity reclassification.
@@ 5.4.1. The antisymmetric matrix
For particles $I<J$ (with initial positions $x_{I}\prec x_{J}$), define the pairwise annihilation weight
@@
For $n=2k$ particles at positions $x_{1}<\cdots<x_{n}$, the total weight of the pairwise coalescence event is
@@
For any ghost sign pattern $(\varepsilon_{1},\ldots,\varepsilon_{k})$, expand $\det(M)$ via the Leibniz formula. With zero survivors, the bijection $\pi\colon\mathcal{A}\to\mathcal{G}$ assigns each actor to one of $k$ ghost pairs (two slots each); the Leibniz terms therefore group naturally by which actors share a ghost pair. The surviving terms are indexed by perfect matchings of $\{1,\ldots,n\}$.
@@
Four particles at $x_{1}<x_{2}<x_{3}<x_{4}$ undergo pairwise coalescence: pair $\{x_{1},x_{2}\}$ and pair $\{x_{3},x_{4}\}$ each merge by time $T$. By parity reclassification, this is equivalent to complete annihilation of all four particles. The $4\times 4$ antisymmetric matrix $A$ has entries $A_{IJ}$ given by (5.1), and the Pfaffian expands as
@@ 5.5.1. Annihilation proves coalescence
The proof uses the annihilation formula to establish a result about coalescence. This interplay between the two models—a payoff of the ghost framework—is made possible by the parity reclassification.
@@ 5.5.3. Connection to gap statistics
The Pfaffian pairwise coalescence formula connects to empty-interval probabilities and voter model duality. In the companion paper [Śni26], this connection is developed further: voter model duality bridges the ghost framework to Pfaffian point process theory [TZ11, GPTZ18], and the Pfaffian formula yields empty-interval probabilities for coalescing random walks.
@@ A.1. Prescribed annihilation
Given $n$ particles at $x_{1}\prec\cdots\prec x_{n}$, fix adjacent particles $x_{m}$ and $x_{m+1}$. The prescribed annihilation weight $Z_{\text{prescribed}}$ is the total weight of all evolutions in which $x_{m}$ and $x_{m+1}$ annihilate (producing a ghost pair at positions $a\prec b$) while the remaining n2n-2 particles reach positions $y_{1}\prec\cdots\prec y_{n-2}$ satisfying $a\prec y_{1}\prec b$, without colliding with each other or with the annihilating pair before annihilation.
@@ A.2. Computation
On the $\mathbb{Z}$ lattice (simple random walk with $\pm 1$ steps), the prescribed annihilation weight can be computed exactly by enumerating all $2^{nt}$ path evolutions and filtering to those satisfying the constraints of Definition˜A.1. For a fixed ordering of the $n$ final positions, distinct final-position tuples provide linear equations in the $n!$ unknowns $c_{\pi}$ of (A.1). The resulting system is solved over $\mathbb{Q}$ in exact arithmetic. (The same technique was used to discover the annihilation formula of the present paper and the coalescence formula of the companion paper [Śni26a]; it succeeds when a formula of the form (A.1) exists.)
@@ Acknowledgments
We thank Maciej Dołęga, Sho Matsumoto, and Karol Życzkowski for stimulating discussions.

### v2

@@ 5.3. The coalescence-to-annihilation map
The classical connection between coalescence and annihilation via parity—assigning independent random labels and propagating by symmetric difference—goes back to Griffeath’s cancellative duality [Gri79] and was made explicit at the particle level by ben-Avraham and Brunet [AB05] (see also Athreya and Swart [AS12] for a rigorous treatment). In the ghost framework, the conversion is deterministic: each entity carries the parity of its multiplicity, and under this cancellative labeling, pairwise coalescence becomes complete annihilation.
@@ 5.4.1. The antisymmetric matrix
For particles $I<J$ (with initial positions $x_{I}\preceq x_{J}$), define the pairwise annihilation weight
@@
For $n=2k$ particles at positions $x_{1}\leq\cdots\leq x_{n}$, the total weight of the pairwise coalescence event is
@@
For any ghost sign pattern $(\varepsilon_{1},\ldots,\varepsilon_{k})$, expand $\det(M)$ via the Leibniz formula. With zero survivors, the bijection $\pi\colon\mathcal{A}\to\mathcal{G}$ assigns each actor to one of $k$ ghost pairs (two slots each); the Leibniz terms therefore group naturally by which actors share a ghost pair. The surviving terms are indexed by perfect matchings of $\{1,\ldots,n\}$. Each perfect matching corresponds to $k!$ candidate bijections (one per numbering of ghost pairs); the $1/k!$ factor in the annihilation formula absorbs this multiplicity.
@@
Four particles at $x_{1}<x_{2}<x_{3}<x_{4}$ undergo pairwise coalescence: pair $\{x_{1},x_{2}\}$ and pair $\{x_{3},x_{4}\}$ each merge by time $T$. By the cancellative labeling, this is equivalent to complete annihilation of all four particles. The $4\times 4$ antisymmetric matrix $A$ has entries $A_{IJ}$ given by (5.1), and the Pfaffian expands as
@@
5.4.2. Random walks with i.i.d. steps
Section˜5.4.1 applies to any spacetime graph satisfying the standing assumptions. A natural special case is $n=2k$ independent random walks on $\mathbb{Z}$ with common step probabilities: at each discrete time step, each particle jumps +1+1 with probability $p$ or 1-1 with probability q=1pq=1-p, independently of all other particles.
Corollary 5.5 (Biased random walk on $\mathbb{Z}$).
For $n=2k$ particles at positions $x_{1}<\cdots<x_{n}$ on $\mathbb{Z}$, each performing independent $\pm 1$ random walks with jump probabilities $(p,q)$ for $T$ steps, the total weight of pairwise coalescence is $\operatorname{Pf}(A)$, where $A$ is the $n\times n$ antisymmetric matrix with entries
(5.3) $A_{IJ}\;=\;2\!\!\sum_{y_{1}<y_{2}}w_{T}(x_{I},y_{2})\,w_{T}(x_{J},y_{1})\;\;+\;\sum_{y}w_{T}(x_{I},y)\,w_{T}(x_{J},y)$
for $I<J$, and the transition weight is
(5.4) $w_{T}(x,y)\;=\;\binom{T}{\frac{T+y-x}{2}}\,p^{\frac{T+y-x}{2}}\,q^{\frac{T-y+x}{2}},$
with $w_{T}(x,y)=0$ unless $\tfrac{T+y-x}{2}$ is a non-negative integer at most $T$. Both sums in (5.3) range over all integers $y$ (respectively pairs $y_{1}<y_{2}$) of the same parity as $x_{I}-T$ and $x_{J}-T$.
Proof.
The spacetime graph is the path graph $\{0,1,\ldots,T\}\times\mathbb{Z}$ with edges $(t,z)\to(t+1,z\pm 1)$ weighted by $p$ and $q$. The path generating function $W(x\to y)=w_{T}(x,y)$ is the binomial transition probability (5.4). Substituting into (5.1) gives (5.3). ∎
The first sum in (5.3) counts crossing path pairs: particle $I$ (starting left) ends strictly right of particle $J$ (starting right). The second sum counts pairs ending at the same position; it vanishes when $x_{J}-x_{I}$ is odd, since the two particles then occupy disjoint sublattices at time $T$. For $p=q=\tfrac{1}{2}$, the transition weight simplifies to $w_{T}(x,y)=\binom{T}{(T+y-x)/2}/2^{T}$.
@@ 5.5.1. Annihilation proves coalescence
The proof uses the annihilation formula to establish a result about coalescence. This interplay between the two models—a payoff of the ghost framework—is made possible by the cancellative labeling. In the physics literature, Masser and ben-Avraham [MA01] showed that the $n$-point correlation functions of annihilation and coalescence are asymptotically identical at large times. Glinyanaya and Fomichov [GF17] proved a central limit theorem for the coalescence cluster count with explicit variance constant $\frac{3-2\sqrt{2}}{\sqrt{\pi t}}$. Our formula makes this connection exact at the level of individual configurations.
@@ 5.5.2. From determinant to Pfaffian
5.5.3. Relation to prior Pfaffian formulas
For Brownian motion, Section˜5.4.1 recovers a known result. Theorem 3 of Tribe and Zaboronski [TZ11] expresses the even product moments of annihilating Brownian motions as Pfaffians of pairwise moments; specializing to the test function $g\equiv 0$ yields the complete extinction probability as $\operatorname{Pf}$ of pairwise extinction probabilities. Garrod, Poplavskyi, Tribe, and Zaboronski [GPTZ18] proved the analogous identity for discrete nearest-neighbor walks (Lemma 7 in [GPTZ18]). Both proofs verify that the Pfaffian satisfies the same system of differential equations as the product moment.
In an earlier, independent direction, Mattera [Mat03] established a bijection between annihilating random walk configurations and perfect matchings (dimers) of a planar spacetime graph, and observed that the remaining particles form a Pfaffian point process. Mattera’s setting is a single specific lattice (simple random walk on $\mathbb{Z}/n\mathbb{Z}$ with synchronous updates), and his Pfaffian is the Kasteleyn Pfaffian of the spacetime graph.
Section˜5.4.1 extends this identity to arbitrary planar weighted DAGs and provides a combinatorial explanation: the Pfaffian appears because annihilating particles pair up in perfect matchings, and marginalizing over ghost positions factorizes across pairs.
@@ 5.5.4. Connection to gap statistics
The Pfaffian pairwise coalescence formula connects to empty-interval probabilities and checkerboard duality. In the companion paper [Śni26a], checkerboard duality bridges the ghost framework to Pfaffian point process theory [TZ11, GPTZ18], and the Pfaffian formula yields empty-interval probabilities for coalescing random walks.
@@ A.1. Prescribed annihilation
Given $n=3$ particles at $x_{1}\prec x_{2}\prec x_{3}$, fix adjacent particles $x_{m}$ and $x_{m+1}$. The prescribed annihilation weight $Z_{\text{prescribed}}$ is the total weight of all evolutions in which $x_{m}$ and $x_{m+1}$ annihilate (producing a ghost pair at positions $a\prec b$) while the remaining particle reaches position $y_{1}$ satisfying $a\prec y_{1}\prec b$, without colliding with the annihilating pair before annihilation.
@@ A.2. Computation
On the $\mathbb{Z}$ lattice (simple random walk with $\pm 1$ steps), the prescribed annihilation weight can be computed exactly by enumerating all $2^{nt}$ path evolutions and filtering to those satisfying the constraints of Definition˜A.1. For a fixed ordering of the $n$ final positions, distinct final-position tuples provide linear equations in the $n!$ unknowns $c_{\pi}$ of (A.1). The resulting system is solved over $\mathbb{Q}$ in exact arithmetic. (The same technique was used to discover the annihilation formula of the present paper and the coalescence formula of the companion paper [Śni26]; it succeeds when a formula of the form (A.1) exists.)
@@ Acknowledgments
We thank Maciej Dołęga, Marcin Lis, Sho Matsumoto, and Karol Życzkowski for stimulating discussions and bibliographic suggestions.